OpenStudy (anonymous):

PLEASE HELP!! I WILL GIVE MEDALS! 6=x+2/3

OpenStudy (adamaero):

subtract the fraction from each side

OpenStudy (anonymous):

im still confused...

OpenStudy (cj49):

18=x+2 x=18-2 x=16

OpenStudy (anonymous):

believe me, so am I...

OpenStudy (anonymous):

6.3-2=x

OpenStudy (cj49):

to get rid of the fraction muliply rhs and lhs by 3

OpenStudy (anonymous):

ok thank u guys so much

OpenStudy (cj49):

yw